# FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### 05 Dec 2023 v1

@@ B.2 Prompt Creation
There are two forms of sessions, structured input sessions and free-form input sessions. Most participants completed the same number of structured input sessions and free-input sessions. The number of sessions completed by each participant are shown in Tab. S1. By default, each participant is expected to complete 9 structured input sessions and 9 free-form input sessions (for the first 13 out of the 33 participants chronologically) or to complete 10 structured input sessions and 10 free-form input sessions (for the last 20 out of the 33 participants chronologically). However, the time consumption highly varies for each participant because of numerous reasons (e.g. Internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
User ID Number of Structured Input Sessions Number of Free-Form Input Sessions Number of Valid Images
@@
Table S1: Information of the FERGI dataset.
B.2.1 Structured Input
@@
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs. S3 and S4.
B.2.2 Free-Form Input
@@
B.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
B.3.1 Image Annotation

### 21 Mar 2024 v2

@@ A.2 Prompt Creation
There are two forms of sessions, structured input (SI) sessions and free-form input (FFI) sessions. Most participants completed the same number of structured input sessions and free-form input sessions. The number of sessions completed by each participant are shown in Tab. S1. By default, each participant is expected to complete 9 structured input sessions and 9 free-form input sessions (for the first 13 out of the 33 participants chronologically) or to complete 10 structured input sessions and 10 free-form input sessions (for the last 20 out of the 33 participants chronologically). However, the time consumption highly varies for each participant because of numerous reasons (e.g. internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
User ID Number of SI Sessions Number of FFI Sessions Number of Valid Images
@@
Table S1: Information of the FERGI dataset. Numbers of structured input (SI) sessions, numbers of free-form input (FFI) sessions, and numbers of valid images for each participant are shown.
A.2.1 Structured Input
@@
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs. S1 and S2.
A.2.2 Free-Form Input
@@
A.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S4 and S5.
A.3.1 Image Annotation